# Federated Continual Instruction Tuning

### v1

@@ A.1 Illustration display of FCIT
Heterogeneous FCIT setting extends the former by allowing each client’s data to come from different datasets in each stage. This requires the global server to mitigate catastrophic forgetting across stages while resolving conflicts among datasets within the same stage. This setting is common in real-world scenarios, such as healthcare systems that need to simultaneously manage multiple disease outbreaks while continuously updating to track their progression and mitigate social risks, and our benchmark effectively address this real-world need, providing a comprehensive evaluation framework for such dynamic challenges.
@@ A.2 Visualization of data heterogeneous
It can be seen that, a smaller $\beta$ leads to greater disparities in data distribution among clients, resulting in more extreme heterogeneity, whereas a larger value of $\beta$ indicates a more uniform distribution.
@@ A.3 Visualization of the dataset
Table B1 and Table B2 present the input images and instruction formats of the 12 selected datasets, which exhibit relatively low average zero-shot performance on the base model LLaVA-v1.5-7B, approximately 30% lower than their fine-tuning performance (See Zero-shot and Centralized MTL in Table 4). This better ensures that these datasets remain unseen or unfamiliar to the base model during training, thereby reducing information leakage.
In our experiments, we design two dataset-level settings: Capability-related and Task-related. The capability-related setting categorizes the 12 datasets into four dimensions: general, math, chart, and other, where each capability consists of a mix of relevant datasets. The task-related setting selects 8 out of the 12 datasets for different stages of continual learning, evaluating how well different approaches mitigate forgetting over long-stage learning setup. The specific data composition of these two settings can be found in Section 3.2.
Appendix B More Details of Experiments
@@
Table B1: Visualization of input images and instruction formats for each dataset in FCIT.
@@
Table B2: Visualization of input images and instruction formats for each dataset in FCIT.
@@ B.3 Ablation study of hyper-parameters
In this section, we conduct more ablation experiments on two key hyper-parameters, the threshold $\tau$ and the temperature coefficient $\varepsilon$, with results shown in Figure B1. It can be seen that both excessively large and small temperature coefficients significantly affect the results. A larger coefficient leads to overly sharp activation assignments, increasing the likelihood of selecting mismatched subspaces, while a smaller coefficient incorporates excessive information from irrelevant subspaces, ultimately degrading model performance. Similarly, an excessively large threshold may filter out knowledge relevant to the same task, while a too-small threshold may misassign identity tokens to other subspaces, both leading to degraded performance.
@@
B.4 Comparison of single Dataset

### v2

@@ 6 Conclusion
Acknowledgments
This work was supported by the National Science and Technology Major Project (2022ZD0116500), National Natural Science Foundation of China (62222609, 62320106010), CAS Project for Young Scientists in Basic Research (YSBR-083), Strategic Priority Research Program of Chinese Academy of Sciences under Grant (XDA0480200), Major Science and Technology Plan Project on the Future Industry Fields of Xiamen City (3502Z20241027), Unveiling and Leading Projects of Xiamen (3502Z20241011) and the InnoHK program.
@@ A.1 Illustration display of FCIT
Heterogeneous FCIT setting extends the former by allowing each client’s data to come from different datasets in each stage. This requires the global server to mitigate catastrophic forgetting across stages while resolving conflicts among datasets within the same stage. This setting is common in real-world scenarios, such as healthcare systems that need to simultaneously manage multiple disease outbreaks while continuously updating to track their progression and mitigate social risks, and our benchmark effectively addresses this real-world need, providing a comprehensive evaluation framework for such dynamic challenges.
@@ A.2 Visualization of data heterogeneous
It can be seen that a smaller $\beta$ leads to greater disparities in data distribution among clients, resulting in more extreme heterogeneity, whereas a larger value of $\beta$ indicates a more uniform distribution.
@@ A.3 Visualization of the dataset
Table A1 and Table A2 present the input images and instruction formats of the 12 selected datasets, which exhibit relatively low average zero-shot performance on the base model LLaVA-v1.5-7B, approximately 30% lower than their fine-tuning performance (See Zero-shot and Centralized MTL in Table 4). This better ensures that these datasets remain unseen or unfamiliar to the base model during training, thereby reducing information leakage.
In our experiments, we design two dataset-level settings: Capability-related and Task-related. The capability-related setting categorizes the 12 datasets into four dimensions: general, math, chart, and other, where each capability consists of a mix of relevant datasets. The task-related setting selects 8 out of the 12 datasets for different stages of continual learning, evaluating how well different approaches mitigate forgetting over a long-stage learning setup. The specific data composition of these two settings can be found in Section 3.2.
@@
Table A1: Visualization of input images and instruction formats for each dataset in FCIT.
@@
Table A2: Visualization of input images and instruction formats for each dataset in FCIT.
Appendix B More Details of Experiments
@@ B.3 Ablation study of hyper-parameters
In this section, we conduct more ablation experiments on two key hyperparameters, the threshold $\tau$ and the temperature coefficient $\varepsilon$, with results shown in Figure B1. It can be seen that both excessively large and small temperature coefficients significantly affect the results. A larger coefficient leads to overly sharp activation assignments, increasing the likelihood of selecting mismatched subspaces, while a smaller coefficient incorporates excessive information from irrelevant subspaces, ultimately degrading model performance. Similarly, an excessively large threshold may filter out knowledge relevant to the same task, while a too-small threshold may misassign identity tokens to other subspaces, both leading to degraded performance.
@@
Threshold $\tau$leq 0.5 0.6 0.7 0.75 0.8 0.9 0.95
Num. of Subspace 1 3 6 7 8 8 10
Table B3: The effect of threshold selection on the number of subspaces in Hom-FCIT setting (task-related, $\beta=1.0$).
We further investigate the effect of threshold selection on the number of subspaces formed. As shown in Table B3, a threshold that is too low fails to effectively separate task-specific knowledge, leading to subspace aggregation across tasks and potential knowledge conflicts. Conversely, an overly high threshold may split knowledge that belongs to the same subspace, resulting in redundant branches and degraded model performance.
B.4 Comparison on a single dataset
@@
B.5 Evaluation of zero-shot capability and general benchmark performance
In addition to mitigating forgetting, we also evaluate the generalization ability of different methods on unseen tasks, as well as their performance on a general LMM benchmark. As shown in Table B4, our method improves zero-shot transfer performance on tasks without downstream supervision. Moreover, it minimizes negative transfer effects on the general LMM benchmark, demonstrating strong robustness and adaptability.
Hom-FCIT $\beta=1.0$ Zero-shot capability General LMM benchmark
Task2 Task3 Task4 Avg MME POPE MMBench SEED
LLaVA-1.5 46.97 17.05 27.13 30.38 1476.9 86.4 66.1 60.1
O-LoRA 41.89 16.04 26.88 28.27 1354.6 80.7 59.4 59.0
M-LoRA 45.33 15.90 31.26 30.83 1376.8 79.8 60.1 59.5
MoELoRA 43.16 16.86 29.09 29.70 1358.3 82.8 60.5 59.2
DISCO 46.14 17.00 36.09 33.08 1436.6 83.9 62.4 60.1
Table B4: Zero-shot transfer and general LMM benchmark results in the Hom-FCIT setting (capability-related, $\beta=1.0$).
B.6 Efficiency and storage analysis.
Methods Efficiency LoRA memory cost
Speed ($\uparrow$) FLOPs ($\downarrow$) Hom-FCIT $\Delta$ Het-FCIT $\Delta$
Finetune 3.46 it/s 8.55 T 22.2 M 0.16% 22.2 M 0.16%
O-LoRA 3.42 it/s 9.35 T 88.7 M 0.63% 177.0 M 1.26%
MoELoRA 3.36 it/s 9.77 T 93.4 M 0.68% 186.0 M 1.33%
DISCO 3.45 it/s 9.35 T 88.9 M 0.63% 178.0 M 1.26%
Table B5: All efficiency comparisons are conducted under identical conditions. $\Delta$: relative percentage to the backbone model.
As shown in Table B5, our method demonstrates the most substantial performance gains compared to fine-tuning and other baselines, while preserving competitive inference efficiency and requiring only modest additional LoRA storage overhead.
B.7 Detailed Results of DISCO.
DISCO General Other Chart Math
Task1 71.26
Task2 67.72 57.23
Task3 64.47 52.75 50.10
Task4 62.92 53.14 46.43 56.15
Table B6: Results matrix of DISCO in Hom-FCIT setting (capability-related, $\beta=1.0$)
DISCO General Other Chart Math
Task1 69.87 55.51
Task2 66.46 54.75 47.81 51.87
Task3 68.73 56.30 50.00 54.47
Task4 68.01 55.49 54.19 57.88
Table B7: Results matrix of DISCO in Het-FCIT setting (capability-related, $\beta=1.0$)
DISCO ImageNet-R ArxivQA IconQA CLEVR-Math OCRVQA Flickr30k FigureQA super-CLEVR
Task1 86.90
Task2 87.10 93.40
Task3 85.88 93.35 65.47
Task4 84.81 93.79 58.70 60.28
Task5 84.80 93.96 59.14 60.84 63.76
Task6 81.60 93.02 58.62 56.24 37.22 54.57
Task7 81.30 92.45 58.27 56.03 25.82 54.46 43.70
Task8 72.56 92.34 55.59 47.72 35.97 52.49 42.92 50.16
Table B8: Results matrix of DISCO in Hom-FCIT setting (task-related, $\beta=1.0$)
DISCO ImageNet-R ArxivQA IconQA CLEVR-Math OCRVQA Flickr30k FigureQA super-CLEVR
Task1 87.99 58.37 64.76 53.39 41.52
Task2 88.35 54.87 53.68 64.35 53.52 42.08 47.04
Task3 76.46 88.1 61.35 57.22 64.24 55.98 41.85 46.88
Task4 84.45 89.28 61.27 55.33 64.04 55.94 40.75 50.00
Task5 83.90 90.45 63.83 58.23 64.07 53.97 39.92 50.18
Task6 85.46 90.71 64.14 62.53 63.73 53.68 40.23 50.14
Task7 84.75 90.65 68.36 60.28 63.17 42.90 40.38 51.08
Task8 73.95 91.75 68.33 59.55 62.52 55.18 43.55 51.16
Table B9: Results matrix of DISCO in Het-FCIT setting (task-related, $\beta=1.0$)